How many liters of a 4% solution of salt should be added to a 24% solution in order to obtain 1100 liters of a 19% solution?
Answer by richwmiller(17219)

a+b=1100,
0.04*a+0.24*b=0.19*1100
a=1100-b
0.04*(1100-b)+0.24*b=209
44-0.04b+0.24*b=209
0.20*b=165
b=825
a=1100-b
a=275 b=825